### v3

# Covariant formulation of electrodynamics in isotropic media revisited

Abstract
We point out some problematic arguments in a recent paper by Pal (2022 Eur. J. Phys. 43 015204), Franklin’s Comment on it (2024 Eur. J. Phys. 45 028002), and Pal’s Reply (2024 Eur. J. Phys. 45 028001).
1 Introduction
Recent papers published in European Journal of Physics dealing with covariant formulation of electrodynamics in isotropic media [1]-[3] contain some problematic points that can be a source of confusion for novices. We believe that the following analysis of the troublesome points might be helpful to the student of relativistic electrodynamics. Section 2 contains a brief review of some basic results of covariant electrodynamics in the vacuum. In Section 3 we discuss Pal’s original paper [1] and point out some mistakes in Franklin’s Comment on it [2] and Pal’s Reply [3].
The present paper was recently rejected based on an extensive report by a Board Member of European Journal of Physics. The report is reproduced verbatim in Appendix as an amusing illustration of how subtle the task of a referee can be.
2 Some basic results of covariant electrodynamics in the vacuum
For the convenience of the reader, in this Section we recall some basic results of covariant electrodynamics in the vacuum which are relevant for the present discussion. Throughout this paper we use the notation and units (Heaviside–Lorentz) of Refs. [1] and [3].
The electromagnetic field tensor $F^{\mu\nu}$ is defined by
$F^{\mu\nu}=\partial^{\mu}A^{\nu}-\partial^{\nu}A^{\mu}\,,$ (1)
where $A^{\mu}$ is the four-vector for electromagnetic potentials which is the four-component combination of the scalar and vector potentials $\varphi$ and $\vec{A}$
$A^{\mu}=(\varphi,A_{x},A_{y},A_{z})\,,$ (2)
and
$\partial_{\mu}\equiv\frac{\partial}{\partial x^{\mu}}\,,\qquad x^{\mu}=\{ct,x,y,z\}\,;$ (3)
the metric employed is
$\eta_{\mu\nu}=\mbox{diag}(+1,-1,-1,-1)\,,$ (4)
so that
$A_{\mu}=\eta_{\mu\nu}A^{\nu}=(\varphi,-A_{x},-A_{y},-A_{z})\,.$ (5)
Greek indices run from 0 to 3.
The electric and magnetic fields $\vec{E}$ and $\vec{B}$ are expressed in terms of the potentials by
$\vec{E}=-\frac{1}{c}\frac{\partial\vec{A}}{\partial t}-\vec{\nabla}\varphi\,,\qquad\vec{B}=\vec{\nabla}\times\vec{A}\,,$ (6)
so that the homogeneous Maxwell’s equations
×$\vec{\nabla}\times\vec{E}=-\frac{1}{c}\frac{\partial\vec{B}}{\partial t}\,,\qquad\vec{\nabla}\cdot\vec{B}=0\,,$ (7)
are satisfied identically. Employing the tensor $F^{\mu\nu}$ equations (7) are written in a manifestly covariant form as
$\partial^{\alpha}F^{\beta\gamma}+\partial^{\beta}F^{\gamma\alpha}+\partial^{\gamma}F^{\alpha\beta}=0\,.$ (8)
which is an identity.
The inhomogeneous Maxwell’s equations in the vacuum,
$\vec{\nabla}\cdot\vec{E}=\rho\,,\qquad\vec{\nabla}\times\vec{B}=\frac{1}{c}\vec{J}+\frac{1}{c}\frac{\partial\vec{E}}{\partial t}\,,$ (9)
where $\rho$ and $\vec{J}$ are the charge and current density, are written in a manifestly covariant form as [^1] Expressing $F^{\mu\nu}$ in terms of $\vec{E}$ and $\vec{B}$, one has $F^{10}=E_{x}\,,F^{20}=E_{y}\,,F^{30}=E_{z}\,,F^{13}=B_{y}\,,F^{21}=B_{z}\,,F^{32}\equiv B_{x}\,;$ for the corresponding dual tensor $\widetilde{F}^{\mu\nu}\equiv\frac{1}{2}\varepsilon^{\mu\nu\lambda\rho}F_{\lambda\rho}$, where $\varepsilon^{\mu\nu\lambda\rho}$ is the completely antisymmetric Levi–Civita symbol, one has $\widetilde{F}^{01}=B_{x}\,,\widetilde{F}^{02}=B_{y}\,,\widetilde{F}^{03}=B_{z}\,,\widetilde{F}^{13}=E_{y}\,,\widetilde{F}^{21}=E_{z}\,,\widetilde{F}^{32}\equiv E_{x}\,.$ The Levi–Civita symbol $\varepsilon^{\mu\nu\lambda\rho}$ is defined as follows: it equals -1 if the indices form an even permutation of 0, 1, 2, 3; it equals +1 if the indices form an odd permutation of 0, 1, 2, 3, and equals 0 otherwise.
$\partial_{\mu}F^{\mu\nu}=\frac{1}{c}J^{\nu}\,,\qquad J^{\nu}=(\rho c,\vec{J})\,.$ (10)
Expressed in terms of the potentials, eqs. (10) read
$(\partial_{\mu}\partial^{\mu})A^{\nu}-\partial^{\nu}(\partial_{\mu}A^{\mu})=\frac{1}{c}J^{\nu}\,.$ (11)
Choosing the Lorenz gauge, $\partial_{\mu}A^{\mu}=0$, the simplified equations for the potentials are obtained
$\Box A^{\nu}=\frac{1}{c}J^{\nu}\,,\qquad\Box\equiv\partial_{\mu}\partial^{\mu}=\frac{1}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}-\vec{\nabla}^{2}\,.$ (12)
Recall that $\varphi$ and $\vec{A}$ must satisfy the simplified equations (12) but also the Lorenz gauge; the last condition is fulfilled if the continuity equation is satisfied. The potentials $\varphi$ and $\vec{A}$ that satisfy the simplified equations (12) and the Lorenz gauge, satisfy also the inhomogeneous Maxwell’s equations in the vacuum (11).
An analogous remark is also valid if an arbitrary gauge is introduced into eqs. (11): $\varphi$ and $\vec{A}$ must satisfy the corresponding simplified equations but also the gauge chosen; the last condition is fulfilled if the continuity equation is satisfied [9].
Recall that eqs. (11) are covariant only under the assumption that $A^{\mu}$ and $J^{\mu}$ are four-vectors [8]. It should be stressed that $A^{\mu}$ is a four-vector by assumption, regardless of whether a covariant or non-covariant gauge is employed in its calculation in the original inertial frame. Clearly, only choosing a covariant gauge yields the covariant simplified equations for the potentials. But, as $A^{\mu}$ is a true four-vector by assumption, the true Maxwell’s equations (8) and (11) are always covariant, employing a non-covariant or covariant gauge in the original frame.
For example, one can use the Coulomb–gauge four–potential $A^{\mu}_{\mbox{\tiny C}}=(\varphi_{\mbox{\tiny C}},\vec{A}_{\mbox{\tiny C}})$ in a frame $S$ and Lorentz–transform it to another frame $S^{\prime}$ to obtain $A^{\prime\mu}_{\mbox{\tiny C}}=(\varphi^{\prime}_{\mbox{\tiny C}},\vec{A}^{\prime}_{\mbox{\tiny C}})$. Since the Coulomb gauge condition $\vec{\nabla}\cdot\vec{A}_{\mbox{\tiny C}}=0$ is not covariant, $\vec{\nabla}^{\prime}\cdot\vec{A}^{\prime}_{\mbox{\tiny C}}$ does not vanish. But, $\vec{E}$ and $\vec{B}$ are gauge invariant, so one obtains the same $\vec{E}$ and $\vec{B}$ in the original $S$ frame either using the Coulomb gauge or a covariant gauge such as the Lorenz gauge, $\partial_{\mu}A^{\mu}_{\mbox{\tiny L}}=0$, see, e.g., [9]. On the other hand, since the field equations written in terms of the potentials do obey a covariant set of equations, eqs. (8) and (11), one obtains the same $\vec{E}^{\prime}$ and $\vec{B}^{\prime}$ in $S^{\prime}$, regardless of which gauge is used in the $S$ frame (the gauge employed need not be covariant). An instructive example is found in Reference [10].
Finally, we present three four-vectors frequently used in [1].
First, Pal defines a dimensionless velocity four-vector
$u^{\mu}=\gamma\{1,\vec{v}/c\}\,.$ (13)
where $\vec{v}$ is ordinary velocity three-vector and $\gamma\equiv(1-v^{2}/c^{2})^{-1/2}$. He also defines what he calls the electric field four-vector and magnetic field four-vector, $E^{\mu}$ and $B^{\mu}$, respectively, by eqs. (16) of [1]
$E^{\mu}=F^{\mu\nu}u_{\nu}\,.$ (14)
$B^{\mu}=-\frac{1}{2}\varepsilon^{\mu\nu\lambda\rho}u_{\nu}F_{\lambda\rho}=-\widetilde{F}^{\mu\nu}u_{\nu}\,.$ (15)
Using eqs. (13) - (15) one obtains
$E^{\mu}=\gamma(\vec{E}\cdot\vec{v}/c,\vec{E}+\vec{v}\times\vec{B}/c)\,,$ (16)
$B^{\mu}=\gamma(\vec{B}\cdot\vec{v}/c,\vec{B}-\vec{v}\times\vec{E}/c)\,.$ (17)
One can verify that the electromagnetic field tensor $F^{\mu\nu}$ can be reconstructed from $E^{\mu}$ and $B^{\mu}$ as follows (see. e.g., [5])
$F^{\mu\nu}=E^{\mu}u^{\nu}-E^{\nu}u^{\mu}+\varepsilon^{\mu\nu\lambda\rho}B_{\lambda}u_{\rho}\,.$ (18)
The last equation is essential for Pal’s covariant extension of electrodynamics to isotropic media [1].
3 Covariant electrodynamics in isotropic media revisited
3.1 Comments on Pal’s original paper [1]
In Section 3 of [1], which deals with covariant electrodynamics in a moving isotropic medium, Pal states: “We define the three–vectors $\vec{D}$ and $\vec{H}$ whose sources are the free charges and currents, and write the inhomogeneous Maxwell equation in terms of them. These vectors are assumed to be linearly related to the electric and magnetic fields $\vec{E}$ and $\vec{B}$:
$\vec{D}=\epsilon\vec{E}\,,\qquad\vec{H}=\frac{1}{\mu}\vec{B}\,.\mbox{''}$ ” (19)
However, in the familiar Minkowski’s electrodynamics of moving media, which is the electrodynamics discussed by Pal, $\vec{D}$ and $\vec{H}$ are defined by
$\vec{D}\equiv\vec{E}+\vec{P}\,,\qquad\vec{H}\equiv\vec{B}-\vec{M}\,,$ (20)
where $\vec{P}$ and $\vec{M}$ are the polarization and magnetization of a moving medium, respectively, and the inhomogeneous Maxwell’s equations in the moving medium read
$\vec{\nabla}\cdot\vec{D}=\rho_{\mbox{\tiny free}}\,,\qquad\vec{\nabla}\times\vec{H}=\frac{1}{c}\vec{J}_{\mbox{\tiny free}}+\frac{1}{c}\frac{\partial\vec{D}}{\partial t}\,.$ (21)
Now since
$\vec{\nabla}\cdot\vec{P}=-\rho_{\mbox{\tiny bound}}\,,\qquad\vec{\nabla}\times\vec{M}+\frac{1}{c}\frac{\partial\vec{P}}{\partial t}=\frac{1}{c}\vec{J}_{\mbox{\tiny bound}}\,,$ (22)
the sources of $\vec{D}$ and $\vec{H}$ are both the free and bound charges and currents. (Recall that $\vec{\nabla}\cdot\vec{H}=-\vec{\nabla}\cdot\vec{M}$; also, in electrostatics ×$\vec{\nabla}\times\vec{D}=\vec{\nabla}\times\vec{P}$.)
Neither the physical fields $\vec{P}$ and $\vec{M}$ nor auxiliary fields $\vec{D}$ and $\vec{H}$ [^2] It is obvious that $\vec{D}\equiv\vec{E}+\vec{P}$ as the sum of two distinct physical fields must be an auxiliary quantity, and analogously for $\vec{H}\equiv\vec{B}-\vec{M}$. appear in Pal’s fundamental equations of his electrodynamics of moving isotropic media. [The only exception are eqs. (19) (eqs. (12) of [1]), which are not assumptions but consequences of definitions (20) and empirical relations $\vec{P}=(\epsilon-1)\vec{E}$ and $\vec{M}=(\mu-1)\vec{B}/\mu$, which are valid for an isotropic medium at rest.] What Pal presents as ‘constitutive relations,’ in a moving medium, which (according to Pal) ‘are the definitions ascribed to Minkowski,’ eqs. (24) of [1], are nothing but identities
$\epsilon E^{\mu}=\epsilon F^{\mu\nu}u_{\nu}\,,\qquad H^{\mu}=\frac{1}{\mu}B^{\mu}$ (23)
since Pal defines $E^{\mu}$ as $F^{\mu\nu}u_{\nu}$, and $H^{\mu}$ as $B^{\mu}/\mu$, cf. eqs. (16a) and (22b) of [1].
On the other hand, the correct constitutive equations for a moving isotropic medium, deduced by Minkowski in 1908, read
$\vec{D}+\vec{v}\times\vec{H}/c=\epsilon(\vec{E}+\vec{v}\times\vec{B}/c)\,,\quad\vec{H}-\vec{v}\times\vec{D}/c=(\vec{B}-\vec{v}\times\vec{E}/c)/\mu\,,$ (24)
where $\vec{v}$ is the velocity of the moving medium [4]-[7].
As is well known, defining a rank-2 antisymmetric tensor $G^{\mu\nu}$ by
$G^{10}\equiv D_{x}\,,G^{20}\equiv D_{y}\,,G^{30}\equiv D_{z}\,,G^{13}\equiv H_{y}\,,G^{21}\equiv H_{z}\,,G^{32}\equiv H_{x}\,,$ (25)
makes it possible to write the inhomogeneous Maxwell’s equations (21) in a manifestly covariant form
$\partial_{\mu}G^{\mu\nu}=\frac{1}{c}J^{\nu}_{\mbox{\tiny free}}\,,\qquad J^{\nu}_{\mbox{\tiny free}}=(\rho_{\mbox{\tiny free}}c,\vec{J}_{\mbox{\tiny free}})\,,$ (26)
in the general case of a moving anisotropic medium. Introducing $D^{\mu}\equiv G^{\mu\nu}u_{\nu}=\gamma(\vec{D}\cdot\vec{v}/c,\vec{D}+\vec{v}\times\vec{H}/c)$, where $G^{\mu\nu}$ is given by eq. (25), and employing eq. (16), the first eq. (24) is recast into the explicitly covariant form $D^{\mu}=\epsilon E^{\mu}$, whereas Pal’s original definition of $D^{\mu}$ as $\epsilon E^{\mu}$, eq. (22a) of [1], is clearly not a constitutive equation, contrary to his assertion in [1]. Similarly, introducing the tensor $H^{\mu\nu}$ dual to $G^{\mu\nu}$ given by eq. (25),
$H^{10}\equiv H_{x}\,,H^{20}\equiv H_{y}\,,H^{30}\equiv H_{z}\,,H^{13}\equiv-D_{y}\,,H^{21}\equiv-D_{z}\,,H^{32}\equiv-D_{x}\,,$ (27)
and $H^{\mu}\equiv H^{\mu\nu}u_{\nu}=\gamma(\vec{H}\cdot\vec{v}/c,\vec{H}-\vec{v}\times\vec{D}/c)$, the second eq. (24) is recast into $H^{\mu}=B^{\mu}/\mu$, employing eq. (17). On the other hand, Pal’s original definition of $H^{\mu}$ as $B^{\mu}/\mu$, eq. (22b) of [1], is not a constitutive equation.
The constitutive equations (24) imply
$\vec{D}=\gamma^{2}\Bigg{\{}\epsilon\left[\vec{E}+\frac{\vec{v}}{c}\times\vec{B}-\left(\vec{E}\cdot\frac{\vec{v}}{c}\right)\frac{\vec{v}}{c}\right]-\frac{1}{\mu}\frac{\vec{v}}{c}\times\left(\vec{B}-\frac{\vec{v}}{c}\times\vec{E}\right)\Bigg{\}}\,,$ (28)
$\vec{H}=\gamma^{2}\Bigg{\{}\epsilon\frac{\vec{v}}{c}\times\left(\vec{E}+\frac{\vec{v}}{c}\times\vec{B}\right)+\frac{1}{\mu}\left[\vec{B}-\frac{\vec{v}}{c}\times\vec{E}-\left(\vec{B}\cdot\frac{\vec{v}}{c}\right)\frac{\vec{v}}{c}\right]\Bigg{\}}\,,$ (29)
Using eqs. (28) and (29) and definitions (20) one obtains the constitutive equations for $\vec{P}$ and $\vec{M}$
$\vec{P}=\gamma^{2}\Bigg{\{}(\epsilon-1)\left[\vec{E}+\frac{\vec{v}}{c}\times\vec{B}-\left(\vec{E}\cdot\frac{\vec{v}}{c}\right)\frac{\vec{v}}{c}\right]+\frac{\mu-1}{\mu}\frac{\vec{v}}{c}\times\left(\vec{B}-\frac{\vec{v}}{c}\times\vec{E}\right)\Bigg{\}}\,,$ (30)
$\vec{M}=\gamma^{2}\Bigg{\{}\frac{\mu-1}{\mu}\left[\vec{B}-\frac{\vec{v}}{c}\times\vec{E}-\left(\vec{B}\cdot\frac{\vec{v}}{c}\right)\frac{\vec{v}}{c}\right]-(\epsilon-1)\frac{\vec{v}}{c}\times\left(\vec{E}+\frac{\vec{v}}{c}\times\vec{B}\right)\Bigg{\}}\,.$ (31)
Thus components of the tensor $G^{\mu\nu}$ given by eq. (25) are expressed eventually in terms only of $\vec{E}$, $\vec{B}$, $\vec{v}$, $\epsilon$ and $\mu$, through eqs. (28) and (29). As can be seen, the last result can be expressed succinctly, ‘in its full covariant glory’ by eq. (21) of [1],
$G^{\mu\nu}=\epsilon(E^{\mu}u^{\nu}-E^{\nu}u^{\mu})+\frac{1}{\mu}\varepsilon^{\mu\nu\lambda\rho}B_{\lambda}u_{\rho}\,,\quad E^{\mu}\equiv F^{\mu\kappa}u_{\kappa}\,,\quad B^{\mu}\equiv-\frac{1}{2}\varepsilon^{\mu\nu\lambda\rho}u_{\nu}F_{\lambda\rho}\,.$ (32)
The validity of eq. (32) is not very obvious, since it is based on non-obvious eq. (18) (eq. (19) of [1]). Moreover, some important aspects of Minkowski’s electrodynamics remain hidden in Pal’s approach which has swallowed $\vec{P}$ and $\vec{M}$. [^3] For example, the constitutive equations (24) and definitions (20) imply $\vec{P}=(\epsilon-1)(\vec{E}+\vec{v}\times\vec{B}/c)+\vec{v}\times\vec{M}/c\,,$ Recall that the term $\vec{v}\times\vec{M}/c$ in the above equation is a consequence of a relativistic effect which has no classical analogue, that a moving magnetic dipole is an electric dipole (see, e.g., [6]). It should be stressed, however, that eq. (32) is a perfectly correct and useful consequence of the definition (25) and constitutive equations (24), expressed in a compact form.
3.2 Mistakes in Franklin’s Comment [2] and Pal’s Reply [3]
In his Comment [2] on Pal’s paper [1], Franklin arrives at a conclusion that the foursome $(\varphi,A_{x},A_{y},A_{z})$ is not a four–vector $A^{\mu}$ in a moving isotropic medium. Consequently, $F^{\mu\nu}=\partial^{\mu}A^{\nu}-\partial^{\nu}A^{\mu}$ is not a tensor and thus $G^{\mu\nu}$ is not a tensor in the moving medium. Franklin eventually concludes that “… the covariant extension of electromagnetism to a polarizable medium proposed by Pal fails.”
Franklin’s argument goes as follows.
In the vacuum, the wave equation (12) “shows $A^{\nu}$ to be a relativistic four-vector. Then, $F^{\mu\nu}$ is seen to be a tensor.” However, in an isotropic medium at rest $\varphi$ and $\vec{A}$ satisfy equations
$\epsilon\left(\frac{\epsilon\mu}{c^{2}}\partial^{2}_{t}-\vec{\nabla}^{2}\right)\varphi=\rho_{\mbox{\tiny free}}\,,$ (33)
$\frac{1}{\mu}\left(\frac{\epsilon\mu}{c^{2}}\partial^{2}_{t}-\vec{\nabla}^{2}\right)\vec{A}=\frac{1}{c}\vec{J}_{\mbox{\tiny free}}\,.$ (34)
as follows from eqs. (6), (19) and (21), under the proviso that a gauge condition
$\vec{\nabla}\cdot\vec{A}+\frac{\epsilon\mu}{c}\frac{\partial\varphi}{\partial t}=0\,,$ (35)
has been employed. Now Franklin argues: “The wave operator now acting on $\varphi$ and $\vec{A}$ [in eqs. (33) and (34)] is no longer a Lorentz scalar so the four-component quantity $A^{\nu}=(\varphi,\vec{A})$ is no longer a four-vector.” I understand that Franklin means by this that $\varphi$ and $\vec{A}$ now satisfy non-covariant equations (33) and (34) and hence the foursome $(\varphi,\vec{A})$ cannot be a four-vector.
However, Franklin’s logic is invalid, both in the vacuum and isotropic medium case. It contradicts the well-established argument of Minkowski’s electrodynamics of moving media, which is the electrodynamics discussed in [1].
First, the foursome $(\varphi,A_{x},A_{y},A_{z})$ is a four–vector $A^{\mu}=(\varphi,A_{x},A_{y},A_{z})$ by definition, and only thus $F^{\mu\nu}\equiv\partial^{\mu}A^{\nu}-\partial^{\nu}A^{\mu}$ is a rank–2 tensor. Only thus the homogeneous Maxwell’s equations expressed in terms of the potentials are written in a manifestly covariant form by eqs. (8). (see, e.g., [8]).
Second, in his Reply [3], Pal has fully refuted Franklin’s criticism of paper [1], noting that eqs. (33)-(35) apply in the rest frame of a moving medium, and hence the covariance of these equations should not be judged from the equations themselves. Pal recalls that“the crucial question is whether we can write a covariant set of equations which reduce to equations [(33) and (34)] in the rest frame of the medium.” Introducing the four-velocity $u^{\mu}=\gamma\{1,\vec{v}/c\}$ of a moving medium in the frame considered, and expressing the terms in eqs. (33)-(35) through $u^{\mu}$ in a covariant way, Pal has shown convincingly that eqs. (33)-(35) written in their “full covariant glory” read
$\epsilon((\epsilon\mu-1)u_{\mu}u_{\nu}\partial^{\mu}\partial^{\nu}+\Box)u_{\lambda}A^{\lambda}=\frac{1}{c}u_{\lambda}J_{\mbox{\tiny free}}^{\lambda}$ (36)
$\frac{1}{\mu}((\epsilon\mu-1)u_{\mu}u_{\nu}\partial^{\mu}\partial^{\nu}+\Box)(\eta^{\mu\nu}-u^{\mu}u^{\nu})A_{\nu}=\frac{1}{c}(\eta^{\mu\nu}-u^{\mu}u^{\nu})J_{\mbox{\tiny free}\,\nu}$ (37)
$\partial_{\mu}A^{\mu}=(1-\epsilon\mu)(u\cdot\partial)(u\cdot A)\,,$ (38)
respectively. [^4] Pal also noted correctly that eq. (36) can be obtained directly from eqs. (26) and (32), expressing $E^{\mu}$ and $B^{\mu}$ in eq. (32) through $A^{\mu}$, contracting eq. (26) with $u^{\nu}$ and employing the gauge condition (38), and similarly for eq. (37).
To summarize, Pal has demonstrated in [3] that reference [2] contains a serious error by claiming that the foursome $(\varphi,A_{x},A_{y},A_{z})$ of electromagnetic potentials is not a four-vector in a moving isotropic medium and that, consequently, Pal’s covariant extension of electromagnetism to isotropic media fails.
It should be stressed, however, that the last paragraph of [3] contains an unfortunate mistake. Arguing that $A^{\mu}$ is a four-vector in the context of electrodynamics in the vacuum, Pal asserts: “The inhomogeneous field equations, written in terms of the potentials, are covariant only if the potentials are assumed to satisfy a covariant gauge condition. In non–covariant gauges such as the Coulomb gauge or the axial gauge, the potentials do not obey a covariant set of equations, as can be verified from any textbook on classical electromagnetic theory.” However, this is not true.
As is recalled in the present Section 2, the inhomogeneous Maxwell’s equations written in terms of the potentials, eqs. (11), are always covariant (assuming of course that $A^{\mu}$ and $J^{\mu}$ are four–vectors [8]), regardless of whether a covariant gauge is used or not in the original frame; on the other hand, only choosing a covariant gauge yields the covariant simplified equations for the potentials.
An analogous remark applies to Minkowski’s electrodynamics of moving isotropic media: eqs. (8) and (26) are the corresponding covariant set of equations for the potentials, expressing $G^{\mu\nu}$ by eq. (32), even if a non–covariant gauge is employed. The four–potential $A^{\mu}_{\mbox{\tiny B}}$, calculated in a non-covariant gauge B, and the four–potential $A^{\mu}_{\mbox{\tiny G}}$, calculated in a covariant gauge G, by solving the corresponding simplified equations for the potentials, non-covariant and covariant, respectively, [^5] The key question in Franklin’s argument in [2] is whether the gauge condition (35), valid in the rest frame of an isotropic medium, can be expressed in a covariant form, valid in an arbitrary inertial frame, since the basic equations (6), (19) and (21) already belong to a covariant set of equations, as discussed above. Pal has shown convincingly that eq. (38) is a covariant generalization of eq. (35) to the moving medium and thus the simplified equations for the potentials (33) and (34) must have covariant generalizations too, as Pal has also demonstrated successfully by eqs. (36) and (37), respectively. Clearly, replacing the covariant gauge (35) with a non-covariant one would yield non-covariant simplified equations for the potentials, but $A^{\mu}$ would still be a four-vector which obeys covariant equations (8) and (26). yield the same $\vec{E}$ and $\vec{B}$ fields in the original frame, and also in any other inertial frame, by Lorentz–transforming $A^{\mu}_{\mbox{\tiny B}}$ or $A^{\mu}_{\mbox{\tiny G}}$. Both four–potentials are true four–vectors from the point of view of relativity. But, “the inhomogeneous field equations written in terms of the potentials,” eqs. (26) via eq. (32), are always covariant, whereas the simplified equations for the potentials are covariant only in a covariant gauge.
Acknowledgment
I thank Vladimir Hnizdo for illuminating correspondence.
Appendix
“Board Member report
This paper is a comment to three papers:
[1] Pal P B 2022 Covariant formulation of electrodynamics in isotropic media Eur. J. Phys. 43 015204
[2] Franklin J 2024 Comment on “Covariant formulation of electrodynamics in isotropic media” Eur. J. Phys. 45 028002
[3] Pal P B 2024 Reply to Comment on “Covariant formulation of electrodynamics in isotropic media” Eur. J. Phys. 45 028001
All three papers, as well as the present, deal with relativistic formulation of electrodynamics in material media. Paper [2] shows that paper [1] contains a serious mistake and in fact is not valid. Paper [3] provides a kind of answer to the criticism of paper [2], although it does not prove that this criticism is not valid. In my opinion the whole case should be closed.
The present paper, however, goes back to [1]. The present author writes: “We believe that the following analysis of the troublesome points might be helpful to the student of relativistic electrodynamics”. In my opinion this is not the case, the paper will not be helpful. The present paper correctly points out flaws in [1]. This has been done in [2], so I do not see the need to do it once more.
Let me give a comment on the present paper – the part where the author criticizes [1] by not mentioning polarization nor magnetization. Electromagnetic field is described by two pairs of vectors, (E,B) and (D,H). The first pair satisfies the homogeneous part of Maxwell’s equations, the second pair – the inhomogeneous part. Both pairs of vectors can be combined into antisymmetric tensors of rank 2 to have a manifestly covariant formulation. Electrodynamics is closed by constitutive relations, namely relations between the two pairs of vectors. According to the present paper the fundamental vectors are P and M (polarization and magnetization) rather than D and H. However, the transformation from the pair D,H to the pair P,M is simply a linear transformation, so it is a matter of taste to say which pair is more fundamental.
Another part of the present paper argues that [2] contains a serious error by claiming that the four-potential for the electromagnetic field is not a relativistic vector. Whether there is an error or not depends on the interpretation of some statements in [2]. I understand the argument of [2] that if the constitutive relations are given in the standard way, D=$\epsilon$E, B=$\mu$H, with $\epsilon$ and $\mu$ different from 1, than the four-potential satisfies a non-relativistic equation and hence is not a four-vector. This is true since the relativistic invariance is violated by the constitutive relations. If, however, the constitutive relations are relativistic, e.g. as given by Minkowski and referred to in the present paper, the four-potential is a relativistic vector. This is more or less obvious and I do not consider the remark in the present paper on the invariance of the four-vector as a significant one.
Concluding, the present paper contains correct statements but in my opinion they do not sufficiently contribute to physics nor physics education to warrant publication in EJP.”
In my view, the above report completely misrepresents references [1] - [3] and the present paper.